OpenStudy (mikey):

Medal and Fan Factor completely 2c5 + 44c4 + 242c3. 2c3(c + 11)2 2(c + 11)2 2c3(c + 11)(c − 11) 2c3(c2 + 22c + 121)

OpenStudy (mathmale):

Mikey: a reminder: Correct way to express exponentiation: Factor completely 2c^5 + 44c^4 + 242c^3 Ambiguous (unclear): Factor completely 2c5 + 44c4 + 242c3

OpenStudy (mikey):

so i divide each by 2

OpenStudy (mathmale):

Looking at 2c^5 + 44c^4 + 242c^3, I see right off that 2and c^3 are common to all terms of your expression. Yes, div. by 2, but also div. by c^3.

OpenStudy (mikey):

@AloneS

OpenStudy (mikey):

lol im soo lost

OpenStudy (mikey):

idk how to divide by c^3

OpenStudy (mikey):

think the answer is d tho

OpenStudy (mathmale):

c^3 --- = 1 c^3 c^5 --- = c^2 c^3

OpenStudy (mikey):

so you subtract the exponents

OpenStudy (mathmale):

Looks like you need to look up "rules of exponents" and practice applying them.

OpenStudy (mathmale):

Yes. But I'd prefer you look up the rules, copy them down, use and learn them.

OpenStudy (mathmale):

Just so that you know, I need to get off the 'Net now. But I'll be back on later. keep going. You don't need to wait for me.

OpenStudy (mikey):

okay